OpenStudy (aaalive):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (tkhunny):

Why not test it out? 1 + (-1) = 0 -- So much for addition Go!

OpenStudy (aaalive):

I don't know, thats why I asked

OpenStudy (anonymous):

A, C, D... im not sure about d because some of them are undefined

OpenStudy (katie_rogers101):

Division: \frac{-1}{1}=-1 \\
\frac{1}{-1}=-1 \\ \frac{-1}{-1}=1 \\ \frac{1}{1}=1

OpenStudy (katie_rogers101):

that didn't post our right

OpenStudy (tkhunny):

You don't know what? The definition of "Closed"? Experiment! SEE if they are closed. I tried it under addition and it failed. You try it under the other operations and see hwt happens.

OpenStudy (katie_rogers101):

If you work through all of the problems you can determine that subtraction is not closed either

OpenStudy (katie_rogers101):

so now you are left with 2 options

OpenStudy (tkhunny):

There is nothing undefined in this exercise. Is the result in the set or not?

OpenStudy (anonymous):

nm i was thinking interval notation. the answer is division and multiplication

OpenStudy (katie_rogers101):

actually it is not division and multiplication. It is just division.

OpenStudy (tkhunny):

It's also closed under Multiplication.